Company E decided to increase its share capital by capitalizing its reserves. To this end, the old shareholders will be given free 1 new common share in every 4 old shares. The market price of the common stock prior to the severance right is 20 euro. Calculate the price of the common share after the receipt has been cut.
Answer #1

Market Price of Stock PRIOR ti new share issued = 20 Euros

Therefore, Market Price of 4 shares = 20*4 = 80 Euros

New Market Price AFTER issue = Prior Price of 4 shares/5 = 80/5 = 16 Euros per share